OpenStudy (anonymous):

What is the value of m in the following equation: 10m - (7m + 4) - 5m =16? -10 -6 1 5/2

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

10m - 7m - 4 - 5m =16 3m - 4 - 5m = 16 -2m - 4 = 16 -2m = 20 m = -10

OpenStudy (haseeb96):

10m - 7m - 4 - 5m=16 -2m=16+4 m=-20\2 m=-10

OpenStudy (anonymous):

Thanks @Haseeb96 @Baseballa101 and @jeremy02030203

OpenStudy (anonymous):

No prob ;)